# Optimal Monotone Depth-Three Circuit Lower Bounds for Majority

### v1

@@ 4.5.1.1 Case 1:
4.5.1.2 Case 2:
There is a common element for $\{e,f\}$, $\{g,h\}$, and $\{c,d\}$. Let $e=f=d$.
$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
@@ 4.5.1.2 Case 2:
4.5.2 There is a set containing a pair of elements from $\{a,b\}\times\{c,d\}$  – $(\neg\bigwedge_{j=1}^{1}\mathscr{P}(2d,j))\land\mathscr{P}(2d,2)$
   Assume without loss of generality that $\mathbb{F}$ contains a set of the form $\{a,c,e\}$ where $e$ is distinct from $b$ and $d$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
@@
4.5.3 A pair of elements appears in two sets where the third element is from $\{a,b,c,d\}$  – $(\neg\bigwedge_{j=1}^{2}\mathscr{P}(2d,j))\land\mathscr{P}(2d,3)$
Assume that there are two sets $\{e,f,g\}$ and $\{e,f,h\}$ containing the pair $\{e,f\}$ where $g,h\in\{a,b,c,d\}$. Since each of $\{a,b,c,d\}$ appear at most twice, we know that neither $e$ nor $f$ are in $\{a,b,c,d\}$. There are two cases.
4.5.3.1 $g=a$ and $h=c$:
$a$ $c$ $e$ $f$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
@@ 4.5.3.1 $g=a$ and $h=c$:
4.5.3.2 $g=a$ and $h=b$:
$a$ $b$ $e$ $f$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
@@ 4.5.3.2 $g=a$ and $h=b$:
4.5.4 One of $\{a,b,c,d\}$ appears exactly once  – $(\neg\bigwedge_{j=1}^{3}\mathscr{P}(2d,j))\land\mathscr{P}(2d,4)$
Assume that $a$ is unique.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
@@
4.5.5 There exists two sets $\{a,e,f\}$ and $\{b,g,h\}$ where $\{e,f\}\neq\{g,h\}$  – $(\neg\bigwedge_{j=1}^{4}\mathscr{P}(2d,j))\land\mathscr{P}(2d,5)$
We have that every one of $\{a,b,c,d\}$ appears exactly twice in $\mathbb{F}$. Assume we have sets $\{a,e,f\}$ and $\{b,g,h\}$ where $\{e,f\}\neq\{g,h\}$.
$a$ $b$ $e$ $f$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
11 0 n2n-2 t1t-1 s1s-1 $2d$ $\frac{10}{25}$ $\frac{5}{12}$
0 1 n2n-2 t1t-1 s1s-1 $2d$ $\frac{10}{25}$ $\frac{5}{12}$
@@ 4.6 Proof of the Upper bounds for $\Phi_{3}(s,t)$
We now prove the upper bounds for $3$-set systems which contain three $2$-sets as stated in Theorem 4.1. Let $\Xi$} be the class of $3$-set system with three $2$-sets. Let $\mathbb{F}\in{\Xi}$ be a $3$-set system over $n\geq 3$ elements with threshold $t$ and deficit $s$. Without loss of generality, assume $\mathbb{F}$ contains the $2$-sets $\{a,b\}$, $\{c,d\}$, and $\{e,f\}$ where $a,b,c,d,e$ and $f$ are not necessarily distinct. We regard the $2$-sets as edges of a graph, consider several cases based on the structure of the graph and prove the bounds in each case.
4.6.1 The three 2-sets form a path of length 3
Consider the case where $F$ contains three $2$-sets that form a path of length $3$. Let $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$ be the $2$-sets that form a path (of length $3$). We cover the family of such set systems by the following sub families.
$\mathscr{P}(3,1)$:
At least one of $\{b,c\}$ is in no other set than the sets $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$.
$\mathscr{P}(3,2)$:
Both $b$ and $c$ appear in some other set other than the sets $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$
Fact 4.9.
Either $\mathscr{P}(3,1)$ or $\mathscr{P}(3,2)$ holds.
4.6.1.1 At least one of $\{b,c\}$ is in no other set than the sets $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$.  – $\mathscr{P}(3,1)$
Assume without loss of generality that $b$ appears in no other set.
$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
1 0 1 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
0 1 1 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
1 0 1 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
@@ 4.6.1 The three 2-sets form a path of length 3
4.6.1.2 Both $b$ and $c$ appear in some other set other than the sets $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$  – $\mathscr{P}(3,2)$
Assume without loss of generality that $b$ appears in another set which is distinct from the sets in the path. Let $\{b,e,f\}$ such a set. Note the $e$ and $f$ are distinct from $a$ and $c$.
$a$ $b$ $c$ $d$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
1 $\{c,d\}$ remains n1n-1 t1t-1 s2s-2 11 $\frac{5}{7}$ $\frac{2}{3}$
1 0 1 $\{e,f\}$ remains n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$ $\frac{5}{18}$
Total 11 $\frac{17}{18}$
4.6.2 The three 2-sets form a triangle
Let $\mathbb{F}$ be a set system with three 2-sets that form a triangle (cycle of length 3) configuration. Let $\Xi$} be the class of such 3-set systems. We cover $\Xi$} based on the following properties.
$\mathscr{P}(3,1)$:
At least one of $\{a,b,c\}$ is is no other set outside the triangle.
$\mathscr{P}(3,2)$:
At least one of $\{a,b,c\}$ is in exactly one other set outside the triangle, and all are in at least one other set.
$\mathscr{P}(3,3)$:
Each of $\{a,b,c\}$ are in at least two other sets outside the triangle.
Lemma 4.10.
One of $\mathscr{P}(3,1),\mathscr{P}(3,2),\mathscr{P}(3,3)$ always holds for every set system in $\Xi$}.
Proof.
In this case, by $\neg\mathscr{P}(3,1)$ $a,b,c$ are each in at least one clause besides the triangle. By $\neg\mathscr{P}(3,2)$, then, we know none of them is in exactly one clause besides the triangle, as all are in at least one. This means each is in at least two additional clauses, implying $\mathscr{P}(3,3)$. By contradiction, this therefore cannot occur. ∎
4.6.2.1 At least one of $\{a,b,c\}$ is is no other set outside the triangle  – $\mathscr{P}(3,1)$
We must select at least two elements of the triangle, otherwise either one of the clauses would be unsatisfied. Assume without loss of generality that $a$ is in no other set outside of the triangle. In this case, not all the elements in the triangle can be set to 1 simultaneously.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
0 1 1 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
1 0 1 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
1 1 0 n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
Total 11 11
4.6.2.2 At least one of $\{a,b,c\}$ is in exactly one other set outside the triangle, and all are in at least one other set  – $\mathscr{P}(3,2)$
Assume without loss of generality that $a$ is in exactly in one set $\{a,d,e\}$ outside the triangle where $d$ and $e$ are distinct from $b$ and $c$. Note that this set must not contain $b$ or $c$.
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
0 1 1 n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$ $\frac{5}{18}$
1 0 1 n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$ $\frac{5}{18}$
@@ 4.6.2.1 At least one of $\{a,b,c\}$ is is no other set outside the triangle  – $\mathscr{P}(3,1)$
4.6.2.3 Each of $\{a,b,c\}$ are in at least two other sets outside the triangle  – $\mathscr{P}(3,3)$
@@
$a$ $b$ $c$ Consequence $n$ $t$ $s$ Type Even $s$ Odd $s$
0 1 1 n3n-3 t2t-2 s3s-3 22 $\frac{5}{21}$ $\frac{25}{108}$
1 0 1 n3n-3 t2t-2 s3s-3 22 $\frac{5}{21}$ $\frac{25}{108}$
@@
4.6.3 There exists a 2-set that is disjoint from the other two 2-sets
Let $\mathbb{F}$ be a set system with three 2-sets that where one of the 2-sets is disjoint from the remaining 2-set. Without loss of generality, let this isolated 2-set be $\{a,b\}$. Let $\{c,d\}$ and $\{e,f\}$ be the other two sets.
We will first deal with the cases when $s$ is odd. Let $\Xi$} be the class of such 3-set systems. We cover $\Xi$} based on the following properties.
$\mathscr{P}(3,1)$:
$s$ is odd and the other two sets $\{c,d\}$ and $\{e,f\}$ overlap.
$\mathscr{P}(3,2)$:
$s$ is odd, $a$ appears in no other set and $\{c,d\}\cap\{e,f\}=\varnothing$.
$\mathscr{P}(3,3)$:
$s$ is odd, $a$ appears in some other 3-set and $\{c,d\}\cap\{e,f\}=\varnothing$.

### v2

@@ 6.5.1 $\mathscr{P}(2d,1)$ – One of $\{a,b,c,d\}$ appears 3 times
6.5.1.2 Case 2: There is a common variable for $\{e,f\}$, $\{g,h\}$, and $\{c,d\}$.
Let $e=f=d$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $d$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
@@
6.5.2 $\mathscr{P}(2d,2)$ – There is a clause containing a pair of variables from $\{a,b\}\times\{c,d\}$
   Assume that $\mathbb{F}\in{\Xi}_{2d}(s,t)$ satisfies $\mathscr{P}(2d,2)$. Assume without loss of generality that $\mathbb{F}$ contains a clause of the form $\{a,c,e\}$ where $e$ is distinct from $b$ and $d$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
@@ 0 $b=1$ n2n-2 t1t-1 s1s-1 $2o$ $\frac{2}{5}$ $\frac{2}{5}$
6.5.3 $\mathscr{P}(2d,3)$ – A pair of variables appears in two clauses where the third variable is from $\{a,b,c,d\}$
Assume that $\mathbb{F}\in{\Xi}_{2d}(s,t)$ satisfies $\mathscr{P}(2d,3)$. Assume that there are two clauses $\{e,f,g\}$ and $\{e,f,h\}$ containing the pair $\{e,f\}$ where $g,h\in\{a,b,c,d\}$. Since each of $\{a,b,c,d\}$ appear at most twice, we know that neither $e$ nor $f$ are in $\{a,b,c,d\}$. There are three cases.
6.5.3.1 Case 1: $g=a$ and $h=c$:
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
@@ 6.5.3.1 Case 1: $g=a$ and $h=c$:
6.5.3.2 Case 2: $g=a$ and $h=b$:
Further assume that one of $\{e,f\}$ does not appear in any other clause.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $e$ $f$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
@@ 6.5.3.2 Case 2: $g=a$ and $h=b$:
6.5.3.3 Case 3: $g=a$ and $h=b$:
Further assume that both $e$ and $f$ appear in some other clause. Assume that there are clauses $\{e,u,v\}$ and $\{f,x,y\}$. Since we also have that the pair $\{e,f\}$ cannot appear in any other clause, it must be $\{u,v,x,y\}$ is disjoint from $\{e,f\}$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $e$ $f$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
11 0 11 n3n-3 t2t-2 s3s-3 11 $\frac{6}{25}$ $\frac{1}{4}$
11 0 0 1 n4n-4 t2t-2 s2s-2 22 $\frac{1}{6}$ $\frac{1}{6}$
0 11 11 n3n-3 t2t-2 s3s-3 11 $\frac{6}{25}$ $\frac{1}{4}$
0 11 0 1 n4n-4 t2t-2 s2s-2 22 $\frac{1}{6}$ $\frac{1}{6}$
11 11 0 0 n4n-4 t2t-2 s2s-2 22 $\frac{1}{6}$ $\frac{1}{6}$
Total $\frac{49}{50}$ 11
6.5.4 $\mathscr{P}(2d,4)$ – One of $\{a,b,c,d\}$ appears exactly once
Assume that $\mathbb{F}\in{\Xi}_{2d}(s,t)$ satisfies $\mathscr{P}(2d,4)$. Assume that $a$ is unique.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
@@ 6.5.4 $\mathscr{P}(2d,4)$ – One of $\{a,b,c,d\}$ appears exactly once
6.5.5 $\mathscr{P}(2d,5)$ – There exists two clauses $\{a,e,f\}$ and $\{b,g,h\}$ where $\{e,f\}\neq\{g,h\}$
Assume that $\mathbb{F}\in{\Xi}_{2d}(s,t)$ satisfies $\mathscr{P}(2d,5)$. We have that every one of $\{a,b,c,d\}$ appears exactly twice in $\mathbb{F}$. Assume we have clauses $\{a,e,f\}$ and $\{b,g,h\}$ where $\{e,f\}\neq\{g,h\}$ (that is, at least one of $e,f$ is neither $g$ nor $h$, but $e,f,g,h$ may be three total variables).
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
11 0 n2n-2 t1t-1 s1s-1 $2d$ $\frac{2}{5}$ $\frac{5}{12}$
0 1 n2n-2 t1t-1 s1s-1 $2d$ $\frac{2}{5}$ $\frac{5}{12}$
@@ 6.6 Proof of the Upper bounds for $\Phi_{3}(s,t)$
We now prove the upper bounds for monotone $3$-CNFs which contain three $2$-clausess as stated in Theorem 4.1. Let $\mathbb{F}\in{\Xi}_{3}(s,t)$ be the class of such monotone $3$-CNFs over $n\geq 3$ variables with threshold $t$ and deficit $s$. Without loss of generality, assume $\mathbb{F}$ contains the $2$-clauses $\{a,b\}$, $\{c,d\}$, and $\{e,f\}$ where $a,b,c,d,e$ and $f$ are not necessarily distinct. We regard the $2$-clauses as edges of a graph, consider several cases based on the structure of the graph and prove the bounds in each case.
6.6.1 The three $2$-clauses form a path of length 3
Consider the case where $\mathbb{F}$ contains three $2$-clauses that form a path of length $3$. Let $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$ be the $2$-clauses that form a path (of length $3$). We cover the family of such monotone $3$-CNFs with the following sub families.
$\mathscr{P}(3\text{-}{p},1)$:
At least one of $\{b,c\}$ is in no other clause than the clauses $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$.
$\mathscr{P}(3\text{-}{p},2)$:
Both $b$ and $c$ appear in some other clause other than the clauses $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$
6.6.1.1 $\mathscr{P}(3\text{-}{p},1)$ – At least one of $\{b,c\}$ is in no other clause than the clauses $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$.
Assume without loss of generality that $b$ does not appear in any other clause.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $c$ $d$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
1 $b=0,c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
0 1 $b=1$ n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
0 0 $b=d=1$ n4n-4 t2t-2 s2s-2 0 $\frac{2}{7}$ $\frac{7}{27}$
Total $\frac{20}{21}$ $\frac{25}{27}$
6.6.1.2 $\mathscr{P}(3\text{-}{p},2)$ – Both $b$ and $c$ appear in some other clause other than the clauses $\{a,b\}$, $\{b,c\}$ and $\{c,d\}$.
Assume without loss of generality that $b$ appears in another clause which is distinct from the clauses in the path. Let $\{b,e,f\}$ such a clause. Note that $e$ and $f$ are distinct from $a$ and $c$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$b$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
1 n1n-1 t1t-1 s2s-2 11 $\frac{5}{7}$ $\frac{2}{3}$
0 $a=c=1$ n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$ $\frac{5}{18}$
Total 11 $\frac{17}{18}$
6.6.2 The three $2$-clauses form a triangle
Consider the case where $\mathbb{F}\in{\Xi}_{3}(s,t)$ contains three $2$-clauses that form a triangle (cycle of length 3) configuration. Let $\{a,b\}$, $\{b,c\}$ and $\{c,a\}$ be the $2$-clauses that form form the triangle. We cover ${\Xi}_{3}(s,t)$ based on the following properties.
$\mathscr{P}(3\text{-}{t},1)$:
At least one of $\{a,b,c\}$ is is no other clause outside the triangle.
$\mathscr{P}(3\text{-}{t},2)$:
At least one of $\{a,b,c\}$ is in exactly one other clause outside the triangle, and all are in at least one other clause.
$\mathscr{P}(3\text{-}{t},3)$:
Each of $\{a,b,c\}$ are in at least two other clauses outside the triangle.
6.6.2.1 $\mathscr{P}(3\text{-}{t},1)$ – At least one of $\{a,b,c\}$ is is no other clause outside the triangle
We must select at least two variables of the triangle, otherwise either one of the clauses would be unsatisfied. Assume without loss of generality that $a$ is in no other clause outside of the triangle. In this case, not all the variables in the triangle can be clause to 1 simultaneously.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
1 0 $c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
1 1 $c=0$ n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
@@ 0 $b=c=1$ n3n-3 t2t-2 s3s-3 0 $\frac{1}{3}$ $\frac{1}{3}$
6.6.2.2 $\mathscr{P}(3\text{-}{t},2)$ – At least one of $\{a,b,c\}$ is in exactly one other clause outside the triangle, and all are in at least one other clause
Assume without loss of generality that $a$ is in exactly one clause $\{a,d,e\}$ outside the triangle where $d$ and $e$ are distinct from $b$ and $c$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$ $\frac{5}{18}$
1 0 $c=1$ n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$ $\frac{5}{18}$
@@ 0 $b=c=1$ n3n-3 t2t-2 s3s-3 11 $\frac{2}{7}$ $\frac{5}{18}$
6.6.2.3 $\mathscr{P}(3\text{-}{t},3)$ – Each of $\{a,b,c\}$ are in at least two other clauses outside the triangle
@@
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $b$ $c$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Even $s$ Odd $s$
0 $b=c=1$ n3n-3 t2t-2 s3s-3 22 $\frac{5}{21}$ $\frac{25}{108}$
1 0 $c=1$ n3n-3 t2t-2 s3s-3 22 $\frac{5}{21}$ $\frac{25}{108}$
@@ 0 $b=c=1$ n3n-3 t2t-2 s3s-3 22 $\frac{5}{21}$ $\frac{25}{108}$
6.6.3 There exists a $2$-clause that is disjoint from the other two $2$-clauses
Let $\mathbb{F}$ be a monotone $3$-CNF with three $2$-clauses that where one of the 2-clauses is disjoint from the remaining $2$-clauses. Let ${\Xi}_{3\text{-}i}(s,t)$ be the class of such CNFs. Without loss of generality, let this isolated $2$-clauses be $\{a,b\}$. Let $\{c,d\}$ and $\{e,f\}$ be the other two $2$-clauses.
We will first deal with the cases when $s$ is odd. We cover such $\mathbb{F}$ with the following properties.
$\mathscr{P}(3\text{-}{io},1)$:
$s$ is odd and the other two clauses $\{c,d\}$ and $\{e,f\}$ overlap.
$\mathscr{P}(3\text{-}{io},2)$:
$s$ is odd, $a$ appears in no other clause and $\{c,d\}\cap\{e,f\}=\varnothing$.
$\mathscr{P}(3\text{-}{io},3)$:
$s$ is odd, $a$ appears in some other 3-clause and $\{c,d\}\cap\{e,f\}=\varnothing$.
6.6.3.1 $\mathscr{P}(3\text{-}{io},1)$$s$ is odd and the other two clauses $\{c,d\}$ and $\{e,f\}$ overlap.
If $\{c,d\}$ and $\{e,f\}$ overlap, the following calculation applies.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Odd $s$
1 n1n-1 t1t-1 s2s-2 $2o$ $\frac{5}{9}$
0 $b=1$ n2n-2 t1t-1 s1s-1 $2o$ $\frac{4}{9}$
Total 11
6.6.3.2 $\mathscr{P}(3\text{-}{io},2)$$s$ is odd, $a$ appears in no other clause and $\{c,d\}\cap\{e,f\}=\varnothing$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Odd $s$
1 $b=0$ n2n-2 t1t-1 s1s-1 $2d$ $\frac{25}{54}$
0 $b=1$ n2n-2 t1t-1 s1s-1 $2d$ $\frac{25}{54}$
Total $\frac{25}{27}$
6.6.3.3 $\mathscr{P}(3\text{-}{io},3)$$s$ is odd, $a$ appears in some other 3-clause and $\{c,d\}\cap\{e,f\}=\varnothing$.
Core variables Forced variables Restricted CNF parameters Inductive bound
$a$ $n^{\prime}$ $t^{\prime}$ $s^{\prime}$ Type Odd $s$
1 n1n-1 t1t-1 s2s-2 $2d$ $\frac{5}{9}$
0 $b=1$ n2n-2 t1t-1 s1s-1 33 $\frac{7}{18}$
Total $\frac{17}{18}$
We now deal with the case that $s$ is even. We cover ${\Xi}_{3\text{-}i}(s,t)$ with the following properties when $s$ is even.
$\mathscr{P}(3\text{-}{ie},1)$:
$s$ is even and either $a$ or $b$ appear in no other clause.
$\mathscr{P}(3\text{-}{ie},2)$:
$s$ is even, either $a$ or $b$ appears in a 3-clause with a variable from $\{c,d,e,f\}$, and that variable is in at most one of $cd$ and $ef$.
$\mathscr{P}(3\text{-}{ie},3)$:
$s$ is even, either $a$ or $b$ appears in a 3-clause with a variable from $\{c,d,e,f\}$, and that variable is in both of $cd$ and $ef$.
$\mathscr{P}(3\text{-}{ie},4)$:
$s$ is even, $|\{c,d,e,f\}|=3$ and there exists clause $\{v_{ab},x,y\}$ where $v_{ab}\in\{a,b\}$ and $\{x,y\}\cap\{c,d,e,f\}=\varnothing$.
$\mathscr{P}(3\text{-}{ie},5)$:
$s$ is even, $c,d,e,f$ are distinct, there exist $v_{ab}\in\{a,b\}$ and $v_{cdef}\in\{c,d,e,f\}$ such that there exist clauses $\{v_{ab},i,j\}$ and $\{v_{cdef},x,y\}$ where $\{x,y\}\neq\{i,j\}$ and none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$.
$\mathscr{P}(3\text{-}{ie},6)$:
$s$ is even, $c,d,e,f$ are distinct, there exists $v_{ab}\in\{a,b\}$ such that there exist clauses $\{v_{ab},i,j\}$, $\{v_{ab},x,y\}$ where none of $a,b,c,d,e,f$ are part of $\{i,j,x,y\}$.